OpenStudy (anonymous):

substitutuon method help me solve please ! need to show my work 4x+y=18 x=y-3

jimthompson5910 (jim_thompson5910):

Start with 4x+y=18 and replace 'x' with y-3 to get 4x+y=18 4(y-3)+y=18 From here, solve for y

jimthompson5910 (jim_thompson5910):

Once you have a solution in terms of y, use it to find x.